# Localization phenomena in the random XXZ spin chain

### v1

@@ Appendix A Localization types and nomenclature
    with parameters $p,m>0$ that do not depend on the choice of the configuration. That is, the normalized eigenfunction $\psi_{n,\omega}$ is exponentially confined near its localization center $k_{n}$, but the control over the confinement is only semi-uniform (it gets worse for localization centers further away from the origin). Unfortunately, for the Anderson model, the primary model for studying single-particle localization phenomena, SULE localization cannot be upgraded to ULE (uniformly localized eigenvectors). ULE does not occur for this model (and indeed for a broad class of random Schrödinger operators) [DJLS].
    A related form of the eigenstate localization is exponential decay of the eigencorrelator (in expectation), already discussed in Section 2. Roughly speaking, eigencorrelator decay implies SULE localization almost surely (see [AW]).
@@
    If $H_{\omega}$ is completely spectrally localized, one can construct an orthonormal eigenbasis for $\ell^{2}(\Lambda)$, that is, there exists a unitary operatoe $U_{\omega}$ that diagonalizes $H_{\omega}$, i.e., $U_{\omega}^{*}H_{\omega}U_{\omega}$ is a diagonal matrix in the standard basis for $\ell^{2}(\Lambda)$. It turns out that for the Anderson model in the strong disorder regime, for any finite volume $\Lambda\subset\mathbb{Z}^{d}$, with high probability one can construct $U_{\omega}$ which is semi-uniformly quasi-local, meaning that the matrix elements $U_{\omega}(x,y)$ of $U_{\omega}$ satisfy the bound $|U_{\omega}(x,y)|\leq C_{\omega}\langle x\rangle^{p}\mathrm{e}^{-m|x-y|}$ for some $p,m>0$ that do not depend on the choice of the configuration. (This follows from the results in [EK].) The existence of such $U_{\omega}$ not only yields a SULE basis, but also allows to label these eigenfunctions according to the spatial position of their localization centers. Such labeling is possible for the Anderson model due to the fact that one can show that with large probability the spectrum of $H_{\omega}$ is level spaced for sufficiently regular distribution of the random potential. Motivated by the concept of LIOM in the many-body context introduced in Section 1.2, we will refer to this form of the eigenstate localization as semi-uniform LIOM localization, As we already indicated, ULE does not occur for the Anderson model, so one can never upgrade a semi-uniform LIOM localization to a uniform LIOM localization in this context.
@@
    Dynamical localization: In quantum mechanics, the dynamics can be given in either Schrödinger or Heisenberg pictures. For single-particle systems, the Schrödinger picture is more common, whereas in the many-body context it is more natural to consider the Heisenberg picture. A typical object of interest in a single-particle system is the spread, due to the dynamics, of an initially localized wave packet. It can be characterized, for example, by the transport exponent $q$ defined by $\left\lVert X\mathrm{e}^{-itH_{\omega}}\delta_{0}\right\rVert\sim\langle t\rangle^{q}$, where $\delta_{0}$ is the standard basis vector for $\ell^{2}(\Lambda)$ located at the origin, and $X$ is a multiplication operator of the form $(X\psi)(x)=x\psi(x)$. The $q=0$ case is then associated with dynamical localization. We note that complete spectral localization does not imply dynamical localization, even if every eigenfunction is exponentially localized [DJLS1]. SULE and the eigencorrelator exponential decay are sufficient conditions to guarantee that $q=0$.
@@
Dynamical localization can be expressed as non-propagation of information: For any observable $\mathcal{O}_{u}$ supported at site $u\in\mathbb{Z}$, $t\in\mathbb{R}$, and $\ell\in\mathbb{N}$ there exists $m>0$ and an observable $\mathcal{O}_{u,\ell,t}$ supported on $[u]_{\ell}$ such that
@@ Appendix B Exponential sums
The lemma is proven by adapting the argument of [BeW, Lemma B.2], who estimate the case $k=1$ of (B.2).
Let $\mathbf{x}\in\mathcal{P}_{N,k}(\mathbb{Z})$, and suppose that ${\bf x}$ has $m=m_{\mathbf{x}}$ clusters (where $m\in\left\{1,\ldots,k\right\}$). Then ${\bf x}=(x_{1},\ldots,x_{N})$, where $x_{1}<\ldots<x_{N}$, and let $x_{j_{1}}<\ldots<x_{j_{2m}}$, where $j_{1}=1$, $j_{2m}=N$, are the end points for its $m$ clusters. (Note that $m$ and $j_{2},,\ldots,j_{2m-1}$ are $\mathbf{x}$-dependent.) Given $\mathbf{y}\in\mathcal{P}_{N}(\mathbb{Z})$ with $\mathbf{y}=(y_{1},\ldots,y_{N})$, where $y_{1}<\ldots<y_{N}$, we set $t_{i}=y_{i}-x_{i}$, $i=1,\ldots,N$. Then the finite sequences $\tau_{q}=(t_{j_{2q-1}},\ldots,t_{j_{2q}})$ are monotone non-decreasing for each $q=1,\ldots,m$. Let $\mathcal{T}_{q}$ denote the collection of such monotone non-decreasing finite sequences $\tau_{q}$. Let $\mathcal{T}^{\left(N\right)}$ denote the collection of all monotone non-decreasing finite sequences $\tau^{\left(N\right)}=(t_{1},\ldots,t_{N})$.
@@
$\sum_{{\mathbf{x}}\in\mathcal{P}_{N,k}(\mathbb{Z})}\mathrm{e}^{-\alpha\left\lvert\mathbf{x}-\mathbf{y}\right\rvert_{1}}=\sum_{{\mathbf{x}}\in\mathcal{P}_{N,k}(\mathbb{Z})}\prod_{q=1}^{m_{\mathbf{x}}}\mathrm{e}^{-\alpha\sum_{j=2q-1}^{2q}\left\lvert x_{j}-y_{j}\right\rvert}{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\leq}\sum_{m=1}^{k}\sum_{\tau_{1}\in\mathcal{T}_{1},\tau_{2}\in\mathcal{T}_{1},\ldots,\tau_{m}\in\mathcal{T}_{m}}\prod_{q=1}^{m}\mathrm{e}^{-\alpha\sum_{t\in\tau_{q}}\left\lvert t\right\rvert}$ (B.4)
$\leq\sum_{m=1}^{k}\sum_{\tau^{\left(N\right)}_{1},\tau^{\left(N\right)}_{2},\ldots,\tau^{\left(N\right)}_{m}\in\mathcal{T}^{\left(N\right)}}\prod_{q=1}^{m}\mathrm{e}^{-\alpha\sum_{t\in\tau^{\left(N\right)}_{q}}\left\lvert t\right\rvert}{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\leq}\sum_{m=1}^{k}\left(\sum_{\tau^{\left(N\right)}\in\mathcal{T}^{\left(N\right)}}\mathrm{e}^{-\alpha\sum_{t\in\tau^{\left(N\right)}}\left\lvert t\right\rvert}\right)^{m}.$
Given $\tau^{\left(N\right)}\in\mathcal{T}^{\left(N\right)}$, since $\tau^{\left(N\right)}$ is monotone nondecreasing there exists an index $10\leq p\leq N+$ 1, such that such that $t_{j}<0$ for $1\leq j\leq p$ and $t_{j}\geq 0$ for ${p}+1\leq j\leq N$. (Note that sets are allowed to be empty). Thus,
@@
where $n_{q}{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}=n_{q}(\mathbf{x})}=j_{2q}-j_{2q-1}$ for $q=1,2,\ldots,m$.
@@
$\mathcal{S}_{{\bf x},r}$ =$=\left\{{\bf y}\in\mathcal{P}_{N}(\mathbb{Z}),\ d_{H}({\bf x},{\bf y})=r\right\},\quad\mathcal{S}_{{\bf x},r}^{\left(m\right)}=\left\{{\bf y}\in\mathcal{P}_{N}^{\left(m\right)}(\mathbb{Z}),\ d_{H}({\bf x},{\bf y})=r\right\}$ (B.11)
@@
Clearly, the second bound follows immediately from the first one by summing over $m$. To obtain the first bound, we note that ${\bf y}\in\mathcal{S}_{{\bf x},r}^{\left(m\right)}$ implies ${\bf y}\subset[{\bf x}]^{\mathbb{Z}}_{r}$, and hence $\mathbf{y}$ is completely determined by the $2m$ points in $[{\bf x}]^{\mathbb{Z}}_{r}$ that are the end points for its $m$ clusters. Since $\left\lvert[{\bf x}]_{r}\right\rvert\leq N+2{k}r$, we deduce that
@@ Appendix C Quasi-locality of the filter function
    For all $K\subset{\Lambda}$ and $K^{\prime}=[K]_{1}^{K_{1}}$ we have $[P_{-}^{K},H]P_{+}^{K^{\prime}}=0$.
@@
We use the following adaptation of [EK22, Lemma B.1] which does not require $\Lambda$ to be finite.
@@
$H$ satisfies the input conditions (i) and (ii) of [EK22, Lemma B.1], so its output (i.e., (C.1)) is valid as well. ∎
@@
Given $t\in\mathbb{R}_{+}$ and $a\in\mathbb{R}$, let $F_{t,a}$ be the $C^{\infty}$ function on $\mathbb{R}$ given in (4.36). Let $S\subsetneq T\subset\Lambda$, $S,T$ finite, where $S\subset K_{1}$ is is connected in $K_{1}$, and let $\ell=\mathrm{d}_{\mathbb{Z}}\left(S,T^{c}\right)-1$. Then, taking $\beta=\frac{\Delta}{5}$, for all $t>1$, $a\in\mathbb{R}$, and $E\in\mathbb{R}$ we have
$\left\lVert P_{-}^{{S}}F_{t,a}(H-E)P_{+}^{{T}}\right\rVert\leq C\left(\mathrm{e}^{-\frac{1}{2}\ell}+\sqrt{t}\,\mathrm{e}^{-\frac{\beta^{2}\ell^{2}}{8t}}\right),$ (C.2)
@@
In particular, taking $t=\frac{\beta^{2}}{8}\ell$, we have ($\ell\in\mathbb{N}^{0}$)
$\left\lVert P_{-}^{{S}}F_{\frac{\beta^{2}}{8}\ell,a}(H-E)P_{+}^{{T}}\right\rVert\leq C\mathrm{e}^{-\frac{1}{2}\ell}.$ (C.3)
@@
$\hat{F}_{t,0,\varepsilon}(\xi)$ $=\tfrac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\left(-i\sqrt{\tfrac{\pi}{2}}\operatorname{sgn}(\xi-s)\right)\left(\tfrac{1}{\sqrt{2\varepsilon}}\mathrm{e}^{-s^{2}/4\varepsilon}-\tfrac{1}{\sqrt{2t}}\mathrm{e}^{-s^{2}/4t}\right)\,\mathrm{d}s$ (C.5)
$=\tfrac{i}{2\sqrt{2}}\left(\int_{\xi}^{\infty}-\int_{-\infty}^{\xi}\right)\left(\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-s^{2}/4\varepsilon}-\tfrac{1}{\sqrt{t}}\mathrm{e}^{-s^{2}/4t}\right)\,\mathrm{d}s.$
@@
$0=\lim_{\xi\to\infty}\hat{F}_{t,0,\varepsilon}(\xi)=-\tfrac{i}{2\sqrt{2}}\int_{-\infty}^{\infty}\mathrm{e}^{i\lambda s}\left(\frac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-s^{2}/4\varepsilon}-\frac{1}{\sqrt{t}}\mathrm{e}^{-s^{2}/4t}\right)\,\mathrm{d}s,$ (C.6)
@@
$\hat{F}_{t,0,\varepsilon}(\xi)$ $=\tfrac{i}{\sqrt{2}}\int_{\xi}^{\infty}\left(\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-s^{2}/4\varepsilon}-\tfrac{1}{\sqrt{t}}\mathrm{e}^{-s^{2}/4t}\right)\,\mathrm{d}s$
$=-\tfrac{i}{\sqrt{2}}\int_{-\infty}^{\xi}\left(\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-s^{2}/4\varepsilon}-\tfrac{1}{\sqrt{t}}\mathrm{e}^{-s^{2}/4t}\right)\,\mathrm{d}s.$
@@
$\left\lvert\hat{F}_{t,0,\varepsilon}(\xi)\right\rvert\leq\tfrac{1}{\sqrt{2}}\int_{\xi}^{\infty}{\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-s^{2}/4\varepsilon}}\,\mathrm{d}s+\tfrac{1}{\sqrt{2}}\int_{\xi}^{\infty}{\tfrac{1}{\sqrt{t}}\mathrm{e}^{-s^{2}/4t}}\,\mathrm{d}s.$ (C.7)
@@
and recalling $\int_{0}^{\infty}\mathrm{e}^{-y^{2}/2}\,\mathrm{d}y=\sqrt{\frac{\pi}{2}}$, we conclude that
@@
$\int_{\xi}^{\infty}\mathrm{e}^{-s^{2}/4b}\,\mathrm{d}s$ $=\sqrt{2b}\int_{\frac{\xi}{\sqrt{2b}}}^{\infty}\mathrm{e}^{-\frac{y^{2}}{2}}\,\mathrm{d}s\leq\sqrt{\tfrac{\pi\mathrm{e}}{2}}\,\sqrt{2b}\,\mathrm{e}^{-\xi^{2}/4b}$ (C.10)
$\leq\sqrt{\pi\mathrm{e}\,b}\,\mathrm{e}^{-\xi^{2}/4b}\leq 3\sqrt{b}\,\mathrm{e}^{-\xi^{2}/4b}\quad\text{for}\quad\xi\geq 0,\ b>0.$
@@
$\left\lvert\hat{F}_{t,0,\varepsilon}(\xi)\right\rvert$ $\leq\tfrac{3}{\sqrt{2}}\mathrm{e}^{-\xi^{2}/4\varepsilon}+\tfrac{3}{\sqrt{2}}\mathrm{e}^{-\xi^{2}/4t}\leq 5\mathrm{e}^{-\xi^{2}/4t},$ (C.11)
@@
[…]
@@
Hence for $0<\varepsilon<t$ and appropriately chosen value for $c$, say $c=\frac{\Delta}{5}$, (C.11) implies, via Stirling’s approximation, that
@@
Using $\left\lvert(F_{t,0}-F_{t,0,\varepsilon})(x)\right\rvert\leq\varepsilon\left\lvert x\right\rvert$, (C.14), restricting to the N-particle sector $\mathcal{H}_{\Lambda}^{\left(N\right)}$, and recalling thatb $H_{N}$ is a bounded operator, we get
@@
$\leq C\left(\mathrm{e}^{-\frac{1}{2}\ell}+\sqrt{t}\,\mathrm{e}^{-\frac{\Delta^{2}\ell^{2}}{200t}}\right)+\varepsilon\left\lVert H_{N}-E\right\rVert,$
@@
$\left\lVert P_{-}^{{S}}F_{t,a}(H_{N}-E)P_{+}^{{T}}\right\rVert\leq C\left(\mathrm{e}^{-\frac{1}{2}\ell}+\sqrt{t}\,\mathrm{e}^{-\frac{\Delta^{2}\ell^{2}}{200t}}\right)\quad\text{for all}\quad N\in\mathbb{N}.$ (C.16)
@@
[…]

### v3

@@ Appendix A Localization types and nomenclature
    with parameters $p,m>0$ that do not depend on the choice of the configuration. That is, the normalized eigenfunction $\psi_{n,\omega}$ is exponentially confined near its localization center $k_{n}$, but the control over the confinement is only semi-uniform (it gets worse for localization centers further away from the origin). Unfortunately, for the Anderson model, the primary model for studying single-particle localization phenomena, SULE localization cannot be upgraded to ULE (uniformly localized eigenvectors). ULE does not occur for this model (and indeed for a broad class of random Schrödinger operators) [10].
    A related form of the eigenstate localization is exponential decay of the eigencorrelator (in expectation), already discussed in Section 2. Roughly speaking, eigencorrelator decay implies SULE localization almost surely (see [4]).
@@
    If $H_{\omega}$ is completely spectrally localized, one can construct an orthonormal eigenbasis for $\ell^{2}(\Lambda)$, that is, there exists a unitary operatoe $U_{\omega}$ that diagonalizes $H_{\omega}$, i.e., $U_{\omega}^{*}H_{\omega}U_{\omega}$ is a diagonal matrix in the standard basis for $\ell^{2}(\Lambda)$. It turns out that for the Anderson model in the strong disorder regime, for any finite volume $\Lambda\subset\mathbb{Z}^{d}$, with high probability one can construct $U_{\omega}$ which is semi-uniformly quasi-local, meaning that the matrix elements $U_{\omega}(x,y)$ of $U_{\omega}$ satisfy the bound $|U_{\omega}(x,y)|\leq C_{\omega}\langle x\rangle^{p}\mathrm{e}^{-m|x-y|}$ for some $p,m>0$ that do not depend on the choice of the configuration. (This follows from the results in [11].) The existence of such $U_{\omega}$ not only yields a SULE basis, but also allows to label these eigenfunctions according to the spatial position of their localization centers. Such labeling is possible for the Anderson model due to the fact that one can show that with large probability the spectrum of $H_{\omega}$ is level spaced for sufficiently regular distribution of the random potential. Motivated by the concept of LIOM in the many-body context introduced in Section 1.2, we will refer to this form of the eigenstate localization as semi-uniform LIOM localization, As we already indicated, ULE does not occur for the Anderson model, so one can never upgrade a semi-uniform LIOM localization to a uniform LIOM localization in this context.
@@
    Dynamical localization: In quantum mechanics, the dynamics can be given in either Schrödinger or Heisenberg pictures. For single-particle systems, the Schrödinger picture is more common, whereas in the many-body context it is more natural to consider the Heisenberg picture. A typical object of interest in a single-particle system is the spread, due to the dynamics, of an initially localized wave packet. It can be characterized, for example, by the transport exponent $q$ defined by $\left\lVert X\mathrm{e}^{-itH_{\omega}}\delta_{0}\right\rVert\sim\langle t\rangle^{q}$, where $\delta_{0}$ is the standard basis vector for $\ell^{2}(\Lambda)$ located at the origin, and $X$ is a multiplication operator of the form $(X\psi)(x)=x\psi(x)$. The $q=0$ case is then associated with dynamical localization. We note that complete spectral localization does not imply dynamical localization, even if every eigenfunction is exponentially localized [22]. SULE and the eigencorrelator exponential decay are sufficient conditions to guarantee that $q=0$.
@@
Dynamical localization can be expressed as non-propagation of information: For any observable $\mathcal{O}_{u}$ supported at site $u\in\mathbb{Z}$, $t\in\mathbb{R}$, and $\ell\in\mathbb{N}$ there exists $m>0$ and an observable $\mathcal{O}_{u,\ell,t}$ supported on $u]_{\ell}$ such that
@@ Appendix B Exponential sums
In this Appendix we prove bounds for the exponential sums encountered throughout the paper. They are used in the context of subsets $\Lambda\subset\mathbb{Z}$ and fixed number of particles $N\in\mathbb{N}$ in conjunction with the basic bounds $d_{1}^{\Lambda}(\mathbf{x},\mathbf{y})\geq d_{1}(\mathbf{x},\mathbf{y})=d^{\mathbb{Z}}_{1}(\mathbf{x},\mathbf{y})$ and $d^{\Lambda}_{H}(\mathbf{x},\mathbf{y})\geq d_{H}(\mathbf{x},\mathbf{y})=d^{\mathbb{Z}}_{H}(\mathbf{x},\mathbf{y})=\left\lvert\mathbf{x}-\mathbf{y}\right\rvert_{1}$ for $\mathbf{x},\mathbf{y}\in\mathcal{P}_{N}(\Lambda)$. Note also that $\mathcal{P}_{N,k}(\Lambda)\subset\mathcal{P}_{N,k}(\mathbb{Z})$ for $k\in\mathbb{N}$.
@@
The lemma is proven by adapting the argument of [7, Lemma B.2], who estimate the case $k=1$ of (B.2).
Let $\mathbf{x}\in\mathcal{P}_{N,k}(\mathbb{Z})$, and suppose that $\bf x$} has $m=m_{\mathbf{x}}$ clusters (where $m\in\left\{1,\ldots,k\right\}$). Then ${\bf x}=(x_{1},\ldots,x_{N})$, where $x_{1}<\ldots<x_{N}$, and let $x_{j_{1}}<\ldots<x_{j_{2m}}$, where $j_{1}=1$, $j_{2m}=N$, be the end points for its $m$ clusters. (Note that $m$ and $j_{2},,\ldots,j_{2m-1}$ are $\mathbf{x}$-dependent.) Given $\mathbf{y}\in\mathcal{P}_{N}(\mathbb{Z})$ with $\mathbf{y}=(y_{1},\ldots,y_{N})$, where $y_{1}<\ldots<y_{N}$, we set $t_{i}=y_{i}-x_{i}$, $i=1,\ldots,N$. Then the finite sequences $\tau_{q}=(t_{j_{2q-1}},\ldots,t_{j_{2q}})$ are monotone non-decreasing for each $q=1,\ldots,m$. Let $\mathcal{T}_{q}$ denote the collection of such monotone non-decreasing finite sequences $\tau_{q}$. Let $\mathcal{T}^{\left(N\right)}$ denote the collection of all monotone non-decreasing finite sequences $\tau^{\left(N\right)}=(t_{1},\ldots,t_{N})$.
@@
$\sum_{{\mathbf{x}}\in\mathcal{P}_{N,k}(\mathbb{Z})}\mathrm{e}^{-\alpha\left\lvert\mathbf{x}-\mathbf{y}\right\rvert_{1}}=\sum_{{\mathbf{x}}\in\mathcal{P}_{N,k}(\mathbb{Z})}\prod_{q=1}^{m_{\mathbf{x}}}\mathrm{e}^{-\alpha\sum_{j=2q-1}^{2q}\left\lvert x_{j}-y_{j}\right\rvert}\leq\sum_{m=1}^{k}\sum_{\tau_{1}\in\mathcal{T}_{1},\tau_{2}\in\mathcal{T}_{1},\ldots,\tau_{m}\in\mathcal{T}_{m}}\prod_{q=1}^{m}\mathrm{e}^{-\alpha\sum_{t\in\tau_{q}}\left\lvert t\right\rvert}$ (B.4)
$\leq\sum_{m=1}^{k}\sum_{\tau^{\left(N\right)}_{1},\tau^{\left(N\right)}_{2},\ldots,\tau^{\left(N\right)}_{m}\in\mathcal{T}^{\left(N\right)}}\prod_{q=1}^{m}\mathrm{e}^{-\alpha\sum_{t\in\tau^{\left(N\right)}_{q}}\left\lvert t\right\rvert}\leq\sum_{m=1}^{k}\left(\sum_{\tau^{\left(N\right)}\in\mathcal{T}^{\left(N\right)}}\mathrm{e}^{-\alpha\sum_{t\in\tau^{\left(N\right)}}\left\lvert t\right\rvert}\right)^{m}.$
Given $\tau^{\left(N\right)}\in\mathcal{T}^{\left(N\right)}$, since $\tau^{\left(N\right)}$ is monotone non-decreasing there exists an index $0\leq p\leq N+1$, such that $t_{j}<0$ for $1\leq j\leq p$ and $t_{j}\geq 0$ for ${p}+1\leq j\leq N$. (Note that sets are allowed to be empty). Thus,
@@
where $n_{q}=n_{q}(\mathbf{x})=j_{2q}-j_{2q-1}$ for $q=1,2,\ldots,m$.
@@
$\mathcal{S}_{{\bf x},r}$ $=\left\{{\bf y}\in\mathcal{P}_{N}(\mathbb{Z}),\ d_{H}({\bf x},{\bf y})=r\right\},\quad\mathcal{S}_{{\bf x},r}^{\left(m\right)}=\left\{{\bf y}\in\mathcal{P}_{N}^{\left(m\right)}(\mathbb{Z}),\ d_{H}({\bf x},{\bf y})=r\right\}$ (B.11)
@@
Clearly, the second bound follows immediately from the first one by summing over $m$. To obtain the first bound, we note that ${\bf y}\in\mathcal{S}_{{\bf x},r}^{\left(m\right)}$ implies ${\bf y}\subset[{\bf x}]^{\mathbb{Z}}_{r}$, and hence $\mathbf{y}$ is completely determined by the $2m$ points in ${\bf x}]^{\mathbb{Z}}_{r}$ that are the end points for its $m$ clusters. Since $\left\lvert[{\bf x}]_{r}\right\rvert\leq N+2{k}r$, we deduce that
@@ Appendix C Quasi-locality of the filter function
    For all $K\subset{\Lambda}$ and $K^{\prime}=[K]_{1}^{K_{1}}$ we have $P_{-}^{K},H]P_{+}^{K^{\prime}}=0$.
@@
We use the following adaptation of [12, Lemma B.1] which does not require $\Lambda$ to be finite.
@@
$H$ satisfies the input conditions (i) and (ii) of [12, Lemma B.1], so its output (i.e., (C.1)) is valid as well. ∎
@@
Given $t\in\mathbb{R}_{+}$ and $a\in\mathbb{R}$, let $F_{t,a}$ be the $C^{\infty}$ function on $\mathbb{R}$ given in (4.36). Let $S\subsetneq T\subset\Lambda$, $S,T$ finite, where $S\subset K_{1}$ is is connected in $K_{1}$, and let $\ell=d_{\mathbb{Z}}\left(S,T^{c}\right)-1$. Then for all $t>1$, $a\in\mathbb{R}$, and $E\in\mathbb{R}$ we have
$\left\lVert P_{-}^{{S}}F_{t,a}(H-E)P_{+}^{{T}}\right\rVert\leq C\left(\mathrm{e}^{-\frac{1}{2}\ell}+\sqrt{t}\,\mathrm{e}^{-\frac{\Delta^{2}\ell^{2}}{100t}}\right),$ (C.2)
@@
In particular, taking $t=\frac{\Delta^{2}}{50}\ell$, we have ($\ell\in\mathbb{N}^{0}$)
$\left\lVert P_{-}^{{S}}F_{\frac{\Delta^{2}}{50}\ell,a}(H-E)P_{+}^{{T}}\right\rVert\leq C\mathrm{e}^{-\frac{1}{2}\ell}.$ (C.3)
@@
$\hat{F}_{t,0,\varepsilon}(\xi)$ $=\tfrac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\left(-i\sqrt{\tfrac{\pi}{2}}\operatorname{sgn}(\xi-s)\right)\left(\tfrac{1}{\sqrt{2\varepsilon}}\mathrm{e}^{-\frac{s^{2}}{4\varepsilon}}-\tfrac{1}{\sqrt{2t}}\mathrm{e}^{-\frac{s^{2}}{4t}}\right)\,\mathrm{d}s$ (C.5)
$=\tfrac{i}{2\sqrt{2}}\left(\int_{\xi}^{\infty}-\int_{-\infty}^{\xi}\right)\left(\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-\frac{s^{2}}{4\varepsilon}}-\tfrac{1}{\sqrt{t}}\mathrm{e}^{-\frac{s^{2}}{4t}}\right)\,\mathrm{d}s.$
@@
$0=\lim_{\xi\to\infty}\hat{F}_{t,0,\varepsilon}(\xi)=-\tfrac{i}{2\sqrt{2}}\int_{-\infty}^{\infty}\mathrm{e}^{i\lambda s}\left(\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-\frac{s^{2}}{4\varepsilon}}-\tfrac{1}{\sqrt{t}}\mathrm{e}^{-\frac{s^{2}}{4t}}\right)\,\mathrm{d}s,$ (C.6)
@@
$\hat{F}_{t,0,\varepsilon}(\xi)$ $=\tfrac{i}{\sqrt{2}}\int_{\xi}^{\infty}\left(\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-\frac{s^{2}}{4\varepsilon}}-\tfrac{1}{\sqrt{t}}\mathrm{e}^{-\frac{s^{2}}{4t}}\right)\,\mathrm{d}s$
$=-\tfrac{i}{\sqrt{2}}\int_{-\infty}^{\xi}\left(\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-\frac{s^{2}}{4\varepsilon}}-\tfrac{1}{\sqrt{t}}\mathrm{e}^{-\frac{s^{2}}{4t}}\right)\,\mathrm{d}s.$
@@
$\left\lvert\hat{F}_{t,0,\varepsilon}(\xi)\right\rvert\leq\tfrac{1}{\sqrt{2}}\int_{\xi}^{\infty}{\tfrac{1}{\sqrt{\varepsilon}}\mathrm{e}^{-\frac{s^{2}}{4\varepsilon}}}\,\mathrm{d}s+\tfrac{1}{\sqrt{2}}\int_{\xi}^{\infty}{\tfrac{1}{\sqrt{t}}\mathrm{e}^{-\frac{s^{2}}{4t}}}\,\mathrm{d}s.$ (C.7)
@@
and recalling $\int_{0}^{\infty}\mathrm{e}^{-\frac{y^{2}}{2}}\,\mathrm{d}y=\sqrt{\frac{\pi}{2}}$, we conclude that
@@
$\int_{\xi}^{\infty}\mathrm{e}^{-\frac{s^{2}}{4b}}\,\mathrm{d}s$ $=\sqrt{2b}\int_{\frac{\xi}{\sqrt{2b}}}^{\infty}\mathrm{e}^{-\frac{y^{2}}{2}}\,\mathrm{d}s\leq\sqrt{\tfrac{\pi\mathrm{e}}{2}}\,\sqrt{2b}\,\mathrm{e}^{-\frac{\xi^{2}}{4b}}$ (C.10)
$\leq\sqrt{\pi\mathrm{e}\,b}\,\mathrm{e}^{-\frac{\xi^{2}}{4b}}\leq 3\sqrt{b}\,\mathrm{e}^{-\frac{\xi^{2}}{4b}}\quad\text{for}\quad\xi\geq 0,\ b>0.$
@@
$\left\lvert\hat{F}_{t,0,\varepsilon}(\xi)\right\rvert$ $\leq\tfrac{3}{\sqrt{2}}\mathrm{e}^{-\frac{\xi^{2}}{4\varepsilon}}+\tfrac{3}{\sqrt{2}}\mathrm{e}^{-\frac{\xi^{2}}{4t}}\leq 5\mathrm{e}^{-\frac{\xi^{2}}{4t}},$ (C.11)
@@
$\left\lVert P_{-}^{A}\ f(H)\,P_{+}^{B}\right\rVert\leq\int_{\mathcal{R}}\left\lVert P_{-}^{A}\ e^{itH}\,P_{+}^{B}\right\rVert\left\lvert\hat{f}(t)\right\rvert dt+\int_{\mathcal{R}^{c}}\left\lvert\hat{f}(t)\right\rvert dt,$ (C.12)
@@
Hence for $0<\varepsilon<t$ and appropriately chosen value for $c$, say $c=\frac{\Delta}{5}$, using Stirling’s approximation and (C.11), we get
@@
Using $\left\lvert(F_{t,0}-F_{t,0,\varepsilon})(x)\right\rvert\leq\varepsilon\left\lvert x\right\rvert$, (C.14), restricting to the N-particle sector $\mathcal{H}_{\Lambda}^{\left(N\right)}$, and recalling that $H_{N}$ is a bounded operator, we get
@@
$\leq C\left(\mathrm{e}^{-\frac{1}{2}\ell}+\sqrt{t}\,\mathrm{e}^{-\frac{\Delta^{2}\ell^{2}}{100t}}\right)+\varepsilon\left\lVert H_{N}-E\right\rVert,$
@@
$\left\lVert P_{-}^{{S}}F_{t,0}(H_{N}-E)P_{+}^{{T}}\right\rVert\leq C\left(\mathrm{e}^{-\frac{1}{2}\ell}+\sqrt{t}\,\mathrm{e}^{-\frac{\Delta^{2}\ell^{2}}{100t}}\right)\quad\text{for all}\quad N\in\mathbb{N}.$ (C.16)
@@
Declarations
  Data availability
We do not analyze or generate any datasets, because our work proceeds within a theoretical and mathematical approach. One can obtain the relevant materials from the references below.
  Funding and/or Conflicts of interests/Competing interests
Alexander Elgart was supported in part by the NSF under grant DMS-2307093.
The authors have no relevant financial or non-financial interests to disclose.
The authors have no competing interests to declare that are relevant to the content of this article.